OpenStudy (anonymous):

The limit represents the derivative of a function f at a number c. Determine f and c.

OpenStudy (anonymous):

lim[(3+h)^2-9]/h as h-->0

OpenStudy (anonymous):

any guesses for this one?

OpenStudy (anonymous):

See is 9 right?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

actually no, \(c=3\)

OpenStudy (anonymous):

\(f(x)=x^2\) \[f(3)=3^2=9\] \[\frac{f(3+h)-f(3)}{h}=\frac{(3+h)^2-9}{h}\]

OpenStudy (anonymous):

Ah, ok thanks. It makes sense when I look at it laid out like that. I was just thinking that since the numerator ends with f(c) (for the definition of derivative formula) that 9 must have been c.